# Tackling GenAI Copyright Issues: Originality Estimation and Generecization

### v4

@@ Abstract
The rapid progress of generative AI technology has sparked significant copyright concerns, leading to numerous lawsuits filed against AI developers. While various techniques for mitigating copyright issues have been studied, significant risks remain. Here, we propose a genericization method that modifies the outputs of a generative model to make them more generic and less likely to infringe copyright. To achieve this, we introduce a metric for quantifying the level of originality of data in a manner that is consistent with the legal framework. This metric can be practically estimated by drawing samples from a generative model, which is then used for the genericization process. As a practical implementation, we introduce PREGen, which combines our genericization method with an existing mitigation technique. Experiments demonstrate that our genericization method successfully modifies the output of a text-to-image generative model so that it produces more generic, copyright-compliant images. Compared to the existing method, PREGen reduces the likelihood of generating copyrighted characters by more than half when the names of copyrighted characters are used as the prompt, dramatically improving the performance. Additionally, while generative models can produce copyrighted characters even when their names are not directly mentioned in the prompt, PREGen almost entirely prevents the generation of such characters in these cases.
@@ 1 Introduction
To mitigate copyright concerns, previous studies have explored methods for reducing the likelihood of producing outputs that resemble copyrighted training data by modifying the training or inference of generative models [vyas2023, wang2024evaluating, he2024fantasticcopyrightedbeastsnot, ma2024datasetbenchmarkcopyrightinfringement]. Although some methods have been found effective in reducing the likelihood of generating copyrighted characters, a significant risk of generating them remains [he2024fantasticcopyrightedbeastsnot, lee2024talkin].
@@
Focusing on the generation of copyrighted characters, which poses particular copyright challenges [sag2023], we demonstrate that PREGen significantly improves the performance of prompt rewriting accompanied by a nagative prompt (we refer to this as the standard method when there is no ambiguity). In particular, experiments using the COPYCAT benchmark [he2024fantasticcopyrightedbeastsnot] show that when copyrighted characters’ names are used as the prompt, PREGen reduces the likelihood of text-to-image generative models generating images of copyrighted characters by more than half compared to the standard method. Furthermore, although generative models can somewhat accidentally generate copyrighted characters even when the prompt appears harmless [he2024fantasticcopyrightedbeastsnot, lee2024talkin], PREGen almost completely eliminates the generation of copyrighted characters in this scenario.
@@ 2 Quantifying originality
AI developers and users might infringe copyright if generative AI produces outputs that resemble existing works. Specifically, outputs that are substantially similar to original works can potentially infringe copyright in them [crs2023]. This means that infringement is found only when the level of similarity reaches some threshold that it is considered substantial. Importantly, copying must be found on protectable elements of the original work [samuelson2015]. Generic elements are usually deemed unprotectible via application of various legal doctrines such as scènes à faire, which denies copyright protection to works or parts of works that are highly typical or commonly found within the specific genre to which the work belongs [lund2009, hacohen2024]. In a similar vein, the higher the level of originality an work possesses, the lower the level of resemblance required to constitute infringement, reflecting the stronger copyright protection for that work [balganesh2012]. In other words, a close copy of a highly generic work may not be considered to infringe copyright, falling short of the substantial similarity threshold, while moderate resemblance may be sufficient to infringe copyright in a highly original work [vermont2012, balganesh2012]. As a consequence of the differential protection based on the level of originality, generic expressions that posses limited originality are less likely to infringe copyright in other works.
@@
The concept of originality, upon which copyright law bases the strength of its protection, can be decomposed into the likelihood of being distinct from other works and the degree of the distinctiveness [vermont2012, byron2006, khong2007]. We propose the following metric for quantifying originality, inspired by the concept of effective originality from the bio-diversity literature [pavoine2021], which captures this intuition.
@@
Aside from the likelihood and degree of being distinct from other works, as captured in our metric, a common-sense notion of originality might encompass such factors as novelty, ingenuity and artistic merit. However, it is generally understood that the copyright law does not directly concern these qualities of the work [gorman2001, vermont2012]. Thus, it is appropriate not to take them into account when determining the strength of copyright protection beyond the extent that is already captured by our metric.
@@
The capacity of generative models to be trained on a massive corpus of preexisting creations and learning the distribution of data can potentially be leveraged to measure the originality of works [tu2021, franceschelli2022, hacohen2024, hacohen2024similarities, haviv2024imageworththousandwords]. In particular, although the true $P(y|x)$ in Equation 1 is not generally available, generative models trained to learn the real-world data distribution can be used as an approximation in practice. For example, to quantify the originality of an image, a text-to-image generative model can be substituted as $P(y|x)$. In this case, $x$ might be thought of as the prompt that sets the distribution of the model.
Observe that the metric is written in the form of expectation. Therefore, although it is difficult to directly calculate Equation 1 in practice, we can obtain the estimate:
@@
where $y_{i}$ are samples drawn from the generative model. This is an unbiased estimator and, by the laws of large numbers, it converges to the true value as we increase the number of samples.
@@ 3 Genericization
Outputs of generative models can be genericized in a model-agnostic way by internally producing $n$ samples, estimating the originality of each sample, and only outputting the one that has the lowest estimated originality. For computational efficiency, the originality estimate of a sample can be cross-computed using other internally produced samples. In other words, the genericization method internally produces $y_{i},i=1,2,\cdots,n$ and outputs $y_{\mathrm{generic}}$ such that
@@
The resulting output is expected to be generic, as opposed to highly unique or original, relative to the prompt. As long as the prompt describes an object that is sufficiently generic, works that are highly similar to the genericized output will likely be generic and thus receive only weak copyright protection, if any, which requires near-exact copying to constitute infringement. By combining this genericization method with prompt rewriting, we can genericize the outputs even when the prompt directly refers to or closely describes copyrighted content (see Section 5).
@@ 4 Proof of concept
Evaluation protocal
We use SDXL [podell2023sdxl], a text-to-image generative model, to obtain originality estimates (Equation 2) of two images of copyrighted characters, Mario from Super Mario Bros and Winnie the Pooh. We used five different prompts (Table 1) for each image as the conditioning $x$, designed to describe the characters at different levels of abstractness or specificity. The specific prompts are designed to capture the essence of the visual aspects of the copyrighted characters, and other prompts were formed by excerpting parts of the specific prompts. We also apply our genericization method to the outputs of the generative model produced with the specific prompts in Table 1.
@@
Abstract “a mustachioed character" “a bear"
Intermediate “a plump, mustachioed character dressed in brown shoes" “a plump, mustachioed character dressed in blue overalls" “a mustachioed character dressed in brown shoes and white gloves" “a chubby, golden-yellow bear" “a chubby bear with a red shirt" “a golden-yellow bear with a red shirt"
Specific “a plump, mustachioed character dressed in a red hat and shirt, blue overalls, brown shoes, and white gloves" “a chubby, golden-yellow bear with a red shirt"
@@
Figure 1: Images of copyrighted characters and characters generated by the generative model. The leftmost column shows the copyrighted images of which originality values were computed. The subsequent columns display sample images generated by the generative model, starting from those generated by the abstract prompts on the left and moving towards those generated with increasingly specific prompts towards the right.
@@ Results
Our metric successfully assigns higher originality values when the generative model is given more abstract prompts and lower originality values with more specific prompts. This reflects the fact that typical generation with more specific prompts, which more closely describe the copyrighted images, is more likely to be highly similar to copyrighted images. In addition, originality estimates for images of copyrighted characters tend to be higher than typical output from the generative model, especially when more abstract prompts are used (Figure 2). However, our originality estimation method can fail when the distribution learned by the generative model is highly distorted (see Appendix C for a case of failure). PREGen discussed in Section 5 addresses this issue by introducing diversity in the prompts used for internally generated samples.
@@
Samples $y_{i}$ with the lowest estimated originality among those selected as $y_{\textrm{generic}}$ as well as samples $y_{i}$ that are the most similar to the copyrighted images are shown in Appendix B. Visual elements that are highly unique to copyrighted characters tend to be absent from genericized sample, which supports the quantitative results that our genericization method effectively makes the output more generic.
@@ 5 Practical implementation and perfomance
We propose an algorithm to mitigate copyright risks of generative models (Algorithm 1), PREGen (Prompt Rwriting-Enhanced Genericization), [^2] The code is available at: https://github.com/hirochok/PREGen. that combines our genericization method involving originality quantification with the standard prompt rewriting method.
@@
Instead of producing a single image using the clean prompt as in the standard prompt rewriting, PREGen first produces an ensemble of images and selects the most generic one as the output, leveraging our originality estimation method. The use of multiple different rewritten prompts in this process, which are similar to each other but not exactly the same, aims to overcome the issue of some generic keywords and phrases having strong association with copyrighted characters [he2024fantasticcopyrightedbeastsnot] (see also Appendix C, where we show that the phrase “Italian plumber" seems to be strongly associated with Mario). Note that rewritten prompts are produced in $b$ batches where $n$ prompts are generated within each batch (Algorithm 1). Within the same batch, all the previously generated prompts are fed into the language model $L$ to generate the next prompt, a set-up that was found to encourage diversity in the rewritten prompts (see the system prompt template in Appendix F).
@@
Input: Prompt $x$; generative model $G$; language model $L$; system prompts $t,\tau$; parameters $b,n$; metric $d$
@@
DETECT measures the similarity of generated images to copyrighted characters by using a detection system, typically a multimodal model such as GPT-4o, to determine whether a specific copyrighted character appears in the generated content. A lower DETECT score indicates that fewer copyrighted characters were generated, which helps in minimizing potential copyright infringement. CONS measures the consistency of the generated image with the user’s intent using VQAScore [lin2024evaluatingtexttovisualgenerationimagetotext]. It evaluates, in particular, whether the key characteristics (e.g. “cartoon mouse" for Mickey Mouse) requested by the user are present in the generated image. A higher CONS score indicates that the generated content aligns better with the user’s expectations, thereby improving user satisfaction. We use the same list of key characteristics used in the previous study [he2024fantasticcopyrightedbeastsnot]. Note that there is a natural trade-off between DETECT and CONS. In an extreme case, a generative model can refuse generation all together or generate a single harmless image regardless of what the prompt says, which would result in a perfect DETECT score. However, the user would likely not be satisfied with such a model. It is crucial to strike the right balance between DETECT and CONS for a generative model to be copyright safe and useful at the same time.
In addition to DETECT and CONS, to aid our analysis, we define p-CONS, a new metric that measures the alignment of the generation with the input prompt provided by the user. p-CONS can be considered as a more fine-grained measure of prompt consistency than the standard CONS. The exact definition of p-CONS and the details of the anlysis involving p-CONS can be found in Appendix D.
@@
“Batman" “An imposing figure draped in a dark, bat-inspired costume. His athletic build is accentuated by a black bodysuit and a cape, with a distinctive emblem of a bat across his chest. His face is obscured by a cowl, leaving only a grim mouth and sharp eyes visible. His utility belt holds an array of gadgets."
“Captain America" “He is a muscular man, often seen in a patriotic uniform consisting of a star-spangled blue torso, red and white striped midsection, and a white star emblem on the chest. His helmet features an ’A’ symbol, and he wields an indestructible round shield of red, white, and blue."
“Judy Hopps" “She is a petite anthropomorphic rabbit with a strong build and gray fur. Her eyes are large and purple, and her ears are long and pink inside. She typically wears a dark blue police uniform, along with a vest and knee pads. Her expressive face constantly reflects her determination and optimism."
“Mario" “A short, plump man sporting a cap and overalls, both in a vibrant shade of blue. He wears a red shirt and gloves, and is always seen with a thick, bushy moustache. His rosy cheeks, bright blue eyes, and brown loafers complete his look."
“Spider-man" “A superhero donning a full-body suit, predominantly red and blue with a web-like design. His mask covers his entire face, featuring white, reflective lenses rimmed with black. His emblem, a black spider, adorns his chest. He is known for his athletic build and acrobatic agility."
Table 2: Sample direct and indirect anchors. Direct anchoring refers to the scenario where the prompt is the name of the copyrighted characters themselves. In the indirect anchoring scenario, the prompt does not mention any copyright character, but describes a character that closely resembles one or contains keywords and phrases that are associated with copyrighted characters. Here, we use approximately 50-word descriptions generated by GPT-4 as the indirect anchors.
@@
We evaluate the performance of PREGen for three different text-to-image generative models, Playground v2.5 [li2024playgroundv25insightsenhancing], Pixart-$\alpha$ [chen2023pixartalphafasttrainingdiffusion]. For comparison purposes, we also conduct the same experiment without any intervention as well as with the standard prompt rewriting method with negative prompting. Experiments were repeated three times for each configuration, and the mean values are reported together with the standard deviations.

### v5

@@ Abstract
The rapid progress of generative AI technology has sparked significant copyright concerns, leading to numerous lawsuits filed against AI developers. While various techniques for mitigating copyright issues have been studied, significant risks remain. Here, we propose a genericization method that modifies the outputs of a generative model to make them more generic and less likely to infringe copyright. To achieve this, we introduce a metric for quantifying the level of originality of data in a manner that is consistent with the legal framework. This metric can be estimated by drawing samples from a generative model, which is then used for the genericization process. As a practical implementation, we introduce PREGen, which combines our genericization method with an existing mitigation technique. Experiments demonstrate that our genericization method successfully modifies the output of a text-to-image generative model so that it produces more generic, copyright-compliant images. Compared to the existing method, PREGen reduces the likelihood of generating copyrighted characters by more than half when the names of copyrighted characters are used as the prompt, dramatically improving the performance. Additionally, while generative models can produce copyrighted characters even when their names are not directly mentioned in the prompt, PREGen almost entirely prevents the generation of such characters in these cases.
@@ 1 Introduction
To mitigate copyright concerns, previous studies have explored methods for reducing the likelihood of producing outputs that resemble copyrighted training data by modifying the training or inference of generative models [vyas2023, wang2024evaluating, he2024fantasticcopyrightedbeastsnot, ma2024datasetbenchmarkcopyrightinfringement]. Although some methods have been found effective in reducing the likelihood of generating outputs resembling copyrighted materials, significant risks of generating them remains [he2024fantasticcopyrightedbeastsnot, lee2024talkin].
@@
Focusing on the generation of copyrighted characters, which poses particular copyright challenges [sag2023], we demonstrate that PREGen significantly improves the performance of prompt rewriting accompanied by a nagative prompt (we refer to this as the standard method when there is no ambiguity). In particular, experiments using the COPYCAT benchmark [he2024fantasticcopyrightedbeastsnot] show that when copyrighted characters’ names are used as the prompt, PREGen reduces the likelihood of text-to-image generative models generating images of copyrighted characters by more than half compared to the standard method. Furthermore, although generative models can inadvertently generate copyrighted characters even with seemingly harmless prompts [he2024fantasticcopyrightedbeastsnot, lee2024talkin], PREGen almost completely eliminates the generation of copyrighted characters in such scenarios.
@@ 2 Quantifying originality
AI developers and users might infringe copyright if generative AI produces outputs that resemble existing works. Specifically, outputs that are substantially similar to original works can potentially infringe copyright in them [crs2023, lee2024talkin]. Consequently, infringement is found only when the significance of similarity reaches some threshold, involving both qunatitative and qualitative aspects, that it is considered substantial [balganesh2012]. Copying must be found on protectable elements of the original work [samuelson2015], and common or standard expressions are usually deemed unprotectible via application of various legal doctrines. For example, scènes à faire denies copyright protection to works or parts of works that are highly typical or commonly found within the specific genre to which the work belongs [lund2009, hacohen2024]. The substantial similarity test operates in such a way that the higher the level of originality a work possesses–in the sense that it involves more creative and expressive qualities [^1] The meaning of originality here is distinct from the basic requirement under copyright law that a work be independently created and display a minimal level of creativity, which qualifies it as an “original work of authorship” eligible for protection.–the thicker its copyright protection. This means that broader range of copying may be considered infringment. Conversely, works with less originality receive thinner protection, meaning fewer types and forms of copying are likely to be considered infringing [balganesh2012, vermont2012].
@@
The concept of originality, which underpins the strength of copyright protection, can be understood as involving both the likelihood of a work being distinct from others and the degree of that distinctiveness [vermont2012, byron2006, khong2007]. The essence of this notion of originality can be captured, at least at a conceptual level, by the following quantity, drawing inspiration from the idea of effective originality in the biodiversity literature [pavoine2021].
@@
Aside from the likelihood and degree of being distinct from other works, as captured in our metric, a common-sense notion of originality might encompass such factors as novelty, ingenuity and artistic merit. However, it is generally understood that the copyright law does not directly concern these qualities of the work [gorman2001, vermont2012]. Thus, in principle, it is appropriate not to take them into account when determining the strength of copyright protection beyond the extent that is already captured by our metric.
@@
The capacity of generative models to be trained on a massive corpus of preexisting creations and learning the distribution of data can potentially be leveraged to measure the originality of works [tu2021, franceschelli2022, hacohen2024, hacohen2024similarities, haviv2024imageworththousandwords]. In particular, although the true $P(y|x)$ in Equation 1 is not generally available, generative models trained to learn the real-world data distribution can be used as an approximation in practice. For example, to quantify the originality of an image, a text-to-image generative model can be substituted as $P(y|x)$. In this case, $x$ might be thought of as the input prompt that sets the distribution of the generation.
Observe that the originality metric (Equation 1) is written in the form of expectation. Therefore, although it is difficult to directly calculate equation in practice, we can obtain the estimate:
@@
where $y_{i}$ are samples drawn from the generative model. This estimator is unbiased and, by the law of large numbers, converges to the expectation as the number of samples increases, thereby linking it to Equation 1, which provides the theoretical foundation for this metric.
@@ 3 Genericization
Outputs of generative models can be genericized in a model-agnostic way by internally producing $n$ samples, estimating the originality of each sample, and only outputting the one that has the lowest estimated originality. For computational efficiency, the originality estimate of a sample can be cross-computed using other internally produced samples. In other words, the genericization method internally produces samples $y_{i},i=1,2,\cdots,n$ and selects the final output $y_{\mathrm{generic}}$ such that
@@
The resulting output is expected to be generic, as opposed to highly unique or original, relative to the prompt. As long as the prompt describes an object that is sufficiently generic, the final output will likely be generic. By combining this genericization method with prompt rewriting, we can genericize the outputs even when the prompt directly refers to or closely describes copyrighted content (see Section 5).
The underlying philosophy of our method is to avoid infringement by generalizing outputs, ensuring they are not overly specific to any protected work. As a consequence of the varying level of copyright protection as embodied in the substantial similarity test, generic expressions–those that use common or widely shared ideas and themes with limited originality–are less likely to infringe on other copyrighted works. This approach is effective with copyrighted characters too, whose unique features make it easier to meet the threshold for substantial similarity [sag2023], while it has broader applicability to other protected works.
@@ 4 Proof of concept
Procedures
We use SDXL [podell2023sdxl], a text-to-image generative model, to obtain originality estimates (Equation 2) of two images of copyrighted characters, Mario from Super Mario Bros and Winnie the Pooh. We used five different prompts (Table 1) for each image as the conditioning $x$, designed to describe the characters at different levels of abstractness or specificity. The specific prompts are designed to capture the essence of the visual aspects of the copyrighted characters, and other prompts were formed by excerpting parts of the specific prompts.
We also test our genericization method by generating multiple samples and selecting the genericized outputs based on Equation 3. This process was applied to the outputs of SDXL created using the specific prompts listed in Table 1.
@@
Abstract “a mustachioed character” “a bear”
Intermediate “a plump, mustachioed character dressed in brown shoes” “a plump, mustachioed character dressed in blue overalls” “a mustachioed character dressed in brown shoes and white gloves” “a chubby, golden-yellow bear” “a chubby bear with a red shirt” “a golden-yellow bear with a red shirt”
Specific “a plump, mustachioed character dressed in a red hat and shirt, blue overalls, brown shoes, and white gloves” “a chubby, golden-yellow bear with a red shirt”
@@
Figure 1: Images of copyrighted characters and characters generated by the generative model. The leftmost column shows the images of copyrighted characters for which originality values were computed. The subsequent columns display sample images generated by the generative model, starting from those generated using the abstract prompts on the left and moving towards those generated with increasingly specific prompts towards the right.
@@ Results
Our metric successfully assigns higher originality values when the generative model is given more abstract prompts and lower originality values for more specific prompts. This reflects the fact that typical generation with more specific prompts, which more closely describe the copyrighted images, is more likely to be highly similar to copyrighted images. In addition, originality estimates for images of copyrighted characters tend to be higher than typical output from the generative model, especially when more abstract prompts are used (Figure 2). However, our originality estimation method can fail when the distribution learned by the generative model is highly distorted (see Appendix C for a case of failure). PREGen discussed in Section 5 addresses this issue by introducing diversity in the prompts used for internally generated samples.
@@
Samples $y_{i}$ with the lowest estimated originality among those selected as $y_{\textrm{generic}}$ as well as samples $y_{i}$ that are the most similar to the copyrighted images are shown in Appendix B. Although the generative model can produce images that share certain highly unique visual elements with copyrighted characters when specific prompts are used, these elements tend to be absent from the genericized samples. This qualitative observation supports that our genericization method effectively makes the output more generic, complementing the quantitative results. Note that, even though our metric does not explicitly dissect the images into protectable and unprotectable elements, selecting the low-originality samples seems to implicitly filter out highly distinctive, protectable features.
@@ 5 Practical implementation and perfomance
We propose an algorithm to mitigate copyright risks of generative models (Algorithm 1), PREGen (Prompt Rwriting-Enhanced Genericization), [^3] The code is available at: https://github.com/hirochok/PREGen. that combines our genericization method involving originality quantification with the standard prompt rewriting method [he2024fantasticcopyrightedbeastsnot].
@@
In sum, instead of producing a single image using the clean prompt as in the standard prompt rewriting, PREGen first produces an ensemble of images and selects the most generic one as the output, leveraging our originality estimation method. The use of multiple different rewritten prompts, which are similar to each other but not exactly the same, in this process aims to overcome the issue of some generic keywords and phrases having strong association with copyrighted characters [he2024fantasticcopyrightedbeastsnot] (see also Appendix C, where we show that the phrase “Italian plumber” seems to be strongly associated with Mario). Note that rewritten prompts are produced in $b$ batches where $n$ prompts are generated within each batch (Algorithm 1). Within the same batch, all the previously generated prompts are fed into the language model to generate the next prompt, a set-up that was found to encourage diversity in the rewritten prompts (see the system prompt template in Appendix F).
@@
Input: Prompt $x$; generative model $G$; language model $L$; system prompts $t,\tau$; integers $b,n$; distance metric $d$
@@
DETECT measures the similarity of generated images to copyrighted characters by using a detection system, typically a multimodal model such as GPT-4o, to determine whether a specific copyrighted character appears in the generated content. A lower DETECT score indicates that fewer copyrighted characters were generated, which helps in minimizing potential copyright infringement. CONS measures the consistency of the generated image with the user’s intent using VQAScore [lin2024evaluatingtexttovisualgenerationimagetotext]. It evaluates, in particular, whether the key characteristics (e.g. “cartoon mouse” for Mickey Mouse) requested by the user are present in the generated image. A higher CONS score indicates that the generated content aligns better with the user’s expectations, thereby improving user satisfaction. We use the same list of key characteristics used in the previous study [he2024fantasticcopyrightedbeastsnot]. Note that there is a natural trade-off between DETECT and CONS. In an extreme case, a generative model can refuse generation all together or generate a single harmless image regardless of what the prompt says, which would result in a perfect DETECT score. However, the user would likely not be satisfied with such a model. It is crucial to strike the right balance between DETECT and CONS for a generative model to be copyright safe and useful at the same time.
In addition to DETECT and CONS, to aid our analysis, we define p-CONS, a new metric that directly measures the alignment of the generation with the input prompt provided by the user. p-CONS can be considered as a more fine-grained measure of prompt consistency than the standard CONS. The exact definition of p-CONS and the details of the anlysis involving p-CONS can be found in Appendix D.
@@
“Batman” “An imposing figure draped in a dark, bat-inspired costume. His athletic build is accentuated by a black bodysuit and a cape, with a distinctive emblem of a bat across his chest. His face is obscured by a cowl, leaving only a grim mouth and sharp eyes visible. His utility belt holds an array of gadgets.”
“Captain America” “He is a muscular man, often seen in a patriotic uniform consisting of a star-spangled blue torso, red and white striped midsection, and a white star emblem on the chest. His helmet features an ‘A’ symbol, and he wields an indestructible round shield of red, white, and blue.”
“Judy Hopps” “She is a petite anthropomorphic rabbit with a strong build and gray fur. Her eyes are large and purple, and her ears are long and pink inside. She typically wears a dark blue police uniform, along with a vest and knee pads. Her expressive face constantly reflects her determination and optimism.”
“Mario” “A short, plump man sporting a cap and overalls, both in a vibrant shade of blue. He wears a red shirt and gloves, and is always seen with a thick, bushy moustache. His rosy cheeks, bright blue eyes, and brown loafers complete his look.”
“Spider-man” “A superhero donning a full-body suit, predominantly red and blue with a web-like design. His mask covers his entire face, featuring white, reflective lenses rimmed with black. His emblem, a black spider, adorns his chest. He is known for his athletic build and acrobatic agility.”
Table 2: Sample direct and indirect anchors. Direct anchoring refers to the scenario where the prompt is the name of the copyrighted characters themselves. In the indirect anchoring scenario, the prompt does not mention any copyright character, but describes a character that closely resembles one. Here, we use approximately 50-word descriptions generated by GPT-4 as the indirect anchors.
@@
We evaluate the performance of PREGen for three different text-to-image generative models, Playground v2.5 [li2024playgroundv25insightsenhancing], Pixart-$\alpha$ [chen2023pixartalphafasttrainingdiffusion] and SDXL. For comparison purposes, we also conduct the same experiment without any intervention as well as with the standard prompt rewriting method with negative prompting. Experiments were repeated three times for each configuration, and the mean values are reported together with the standard deviations.